32. iven the following informa -a[S03]/At, t/min tion, calculate the average rate, between 10- and 40.0 minutes for the loss of so o.0 10.0 20·0 30.0 40.0 50.0 o.124 o.092 0.068 O.050 0.037 0.028 O.0 0.032 O 056 0.074 0.087 0.096 O.01 0-028 0.037 0 044 0.048 a. 2.18 x 10-4 d. 545 b. 1.8 x 10-3 e. 459 c.1.1l x 10-3 33. The following data were measured for ther eaction: 2C12 (g) SOCI2(g) + c120 (g) → so2 (g) + Experiment ISO21M ICl2l) nitial RaterM/s 1 2 3 4 0.400 0.400 0.400 0.200 0.400 0.200 0.800 0.800 0.2918 0.0730 1.1674 0.5837 What is the rate law for the reaction? k[SO2] k[C12]2 D. rate = k[SO2] [C12] d. rate = k[SO2 ] [C12]2 = a. rate = c. rate e. rate = k[SO2 ] 2 [C12] 34. Let the rate of the reaction 2NH3 (g) → N2 (g) + 3H2 (g) be expressed as -A[NH3]/At. Express the rate of the reaction in terms of Δ[H2]/dt.
Answer #1

Ans 32 : b) 1.8 x 10-3

The change in concentration of SO3 in the given time frame is :

= 0.092 - 0.037

= 0.055 M

The change in time is :

= 40 - 10

= 30 min

So the average rate of :

= 0.055 / 30

= 1.8 x 10-3